EVE shift summary, 9/26/99

Alexander Ledovskoy (ledovskoy@fnal.gov)
Mon, 27 Sep 1999 00:15:55 -0500

--------------79100F0C8ED1D8A1D8EA94A4
Content-Type: text/plain; charset=us-ascii
Content-Transfer-Encoding: 7bit

Shift Summary for EVE shift, 09/26/99 (Sun.)

On Shift: Ricardo, Sasha L.

Hours of beam available at all intensities: 8.5h
Hours of beam available at nominal intensity: 8.5h
Hours of E799 data (special runs not included): 8.1h
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time & intensity:

Tape change, KFRENDs, Pipeline resets, small stuff: spill here and
there

========================================================================

Runs ended this shift:

14736 -- BEAM_799 -- 16.8M
(with 1x prescale on Trigger 10)

Tapes brought to FCC this shift:

========================================================================

Accesses during shift:

None.

========================================================================

Overview of shift:

E799 data taking at 9E12 with 10" Be common absorber.
Tried to run at 1E13 but could not handle the amount of data:
KTEV4 causes dead DAQ errors. Returned to 9E12 intensity

========================================================================

Description of Shift:

run spill time comment
------------------------------------------------------------------------

14736

17:15 2 bottles of Argon-Ethane now on
line , east side

22:00 Online split gives LOW disk space
error
Paged bunch of people. Sasha
Glazov helped to
delete files.

368 CsI 479 dropped out of laser
illumination plots.
Reset pipeline, KFRENDs, problem
fixed.
377 Back to data taking

23:36 Tapes are full, end of 14736

14737 23:46 BEAM_799 started.

=====================================================================
Follow up Notes on Previous Shift

========================================================================

Items to Follow up on Next Shift

Monday, DAY : PLEASE:

(1) put the old HP4Si toner cartridges (at least 4 in control room) into

the FNAL recycling system.

(2) get the "BAD" pieces of electronics which litter the area moved over

to PREP for service.

--------------79100F0C8ED1D8A1D8EA94A4
Content-Type: text/html; charset=us-ascii
Content-Transfer-Encoding: 7bit

<!doctype html public "-//w3c//dtd html 4.0 transitional//en">
 
              Shift Summary for EVE shift, 09/26/99 (Sun.)
 

On Shift:    Ricardo, Sasha L.
 

Hours of beam available at all intensities:         8.5h
Hours of beam available at nominal intensity:       8.5h
Hours of E799 data (special runs not included):     8.1h
Hours of special runs (laser scans, muon runs...):  0.0

Accounting of lost time & intensity:

Tape change, KFRENDs, Pipeline resets, small stuff: spill here and there 
 

========================================================================
Runs ended this shift:
 
      14736 -- BEAM_799 -- 16.8M
                (with 1x prescale on Trigger 10)
 

Tapes brought to FCC this shift:
 
========================================================================
Accesses during shift:

    None.

========================================================================
Overview of shift:

    E799 data taking at 9E12 with 10" Be common absorber.
Tried to run at 1E13 but  could not handle the amount of data:
KTEV4 causes dead DAQ errors. Returned to 9E12 intensity

========================================================================
Description of Shift:

run           spill      time      comment
------------------------------------------------------------------------
14736
 

                            17:15      2 bottles of Argon-Ethane now on line , east side
 
                            22:00      Online split gives LOW disk space error
                                       Paged bunch of people. Sasha Glazov helped to
                                       delete  files.

                  368                  CsI 479 dropped out of laser illumination plots.
                                       Reset pipeline, KFRENDs, problem fixed.
                  377                 Back to data taking

                            23:36      Tapes are full,  end of 14736

14737                  23:46       BEAM_799 started.

=====================================================================
Follow up Notes on Previous Shift
 
 

========================================================================
Items to Follow up on Next Shift

Monday, DAY : PLEASE:

(1) put the old HP4Si toner cartridges (at least 4 in control room) into
the FNAL recycling system.

(2) get the "BAD" pieces of electronics which litter the area moved over
to PREP for service.
 
  --------------79100F0C8ED1D8A1D8EA94A4--